OpenStudy (anonymous):

If a is divisible by 5 and b is divisible by 7, which of the following mys be divisible by 35? I. ab II. 7a+5b III. 5a+7b Can have up to 3 answers max and 1 answer min.

OpenStudy (blockcolder):

From the hypotheses, we can say that a=5k and b=7m, where k and m are integers. Let's look at each choice: ab=(5k)(7m)=35(km) 7a+5b=7(5k)+5(7m)=35k+35m=35(k+m) 5a+7b=5(5k)+7(7m)=25k+49m We can then see that ab and 7a+5b are divisible by 35.